OpenStudy (anonymous):

Simplify cos2 Θ(1 + tan2 Θ). cos2 Θ + sin2 Θ cos2 Θ 1 −1

OpenStudy (anonymous):

@ganeshie8

OpenStudy (ipwnbunnies):

Here are some useful trig identites: \[1 + \tan^2 u = \sec^2 u\] \[\sec u = \frac{1}{\cos u}\] The second one also applies if they are raised to some power.

OpenStudy (anonymous):

So the answer would be b right @iPwnBunnies

OpenStudy (ipwnbunnies):

No. What did you do to simpify?

OpenStudy (ipwnbunnies):

\[ = (\cos^2 u)(\sec^2 u)\] \[= \frac{\cos^2 u}{\cos^2 u}\] \[= 1\]

OpenStudy (anonymous):

If sin Θ = negative square root 3 over 2 and π < Θ < 3 pi over 2, what are the values of cos Θ and tan Θ? cos Θ = negative 1 over 2; tan Θ = square root 3 cos Θ = negative 1 over 2; tan Θ = −1 cos Θ = square root 3 over 4; tan Θ = −2 cos Θ = 1 over 2; tan Θ = square root 3 @iPwnBunnies

OpenStudy (anonymous):

I believe it was the third choice?

OpenStudy (anonymous):

@ganeshie8